OpenStudy (anonymous):

(5+√11)(5-√11) Please help i give medals! Answer choices: A.36 B.14 C.25-10√11 D.25-√11

OpenStudy (tyteen4a03):

Hint: \(x^2 - y^2 = (x+y)(x-y)\)

OpenStudy (anonymous):

What formula is that called? It's not the quadratic one is it?

OpenStudy (tyteen4a03):

It's the difference of two squares identity.

OpenStudy (anonymous):

ugghh I'm not sure how to do that

OpenStudy (tyteen4a03):

Just observe the formula and fill in the blanks. What would x and y stand for in the case of \((5+\sqrt{11})(5-\sqrt{11})\)?

OpenStudy (anonymous):

x would be 5 and y would be sqrt11, im not sure though. (sorry it took me so long to rrespond my computer shut down for an update and i had to switch to another computer.

OpenStudy (tyteen4a03):

Correct. Now apply that to \(x^2 - y^2\)

OpenStudy (anonymous):

X would be 25 i think and if you squared a radical sign it should cancel leaving 11.

OpenStudy (anonymous):

im not sure though

OpenStudy (tyteen4a03):

Correct.

OpenStudy (tyteen4a03):

However, x^2 = 25, not x = 25.

OpenStudy (anonymous):

so thats the answer?

OpenStudy (anonymous):

D

OpenStudy (tyteen4a03):

No. x = 5, y = sqrt(11) x^2 = 25, y = 11 We know that x^2 - y^2 = (x+y)(x-y) So what is x^2 - y^2?

OpenStudy (anonymous):

25-11=14

OpenStudy (anonymous):

or do you have to square the 11 to?

OpenStudy (tyteen4a03):

No. It's asking y^2, not y.

OpenStudy (anonymous):

so 121?

OpenStudy (anonymous):

d because 11 cant be broken down and 5*5 is 25

OpenStudy (tyteen4a03):

@callie2240 Please read my solution. @csimmons3 You must be very clear of what you are doing here. Back to first step: You defined x = 5 and y = sqrt(11), this doesn't change for the whole progress. The answer is x^2 - y^2, so you find x^2 and y^2 and substitute them into the equation. You don't magically transform y^2 into y - they stay two different (but related) numbers.

OpenStudy (tyteen4a03):

*substitute into the expression

OpenStudy (anonymous):

i dont see a solution???

OpenStudy (anonymous):

I'm lost

OpenStudy (anonymous):

Ohhhh is it B?

OpenStudy (tyteen4a03):

Correct. Next time you do math, in order to make yourself very clear about what you're doing, write variables and their values out. This way, you won't be confused. @callie2240 See all my posts for details.

OpenStudy (anonymous):

Thanks for all the help :)

OpenStudy (anonymous):

i saw thanks

OpenStudy (anonymous):

it would be B